Find the distance between the points with the given coordinates (4,-3) and (-4, 3)
Answer by Jstrasner(112)

Hey,
For this one, we need to use the pythagorean theorem
which is a^2 + b^2 = c^2
We need to find the two base sides of the triangle that these coordinates can make:
the length of one side is 3-(-3) = 6
The length of the other side is 4-(-4) = 8
6^2 + 8^2 = c^2
36 + 64 = 100
The square root of 100 is 10
10 is the distance
I hope this helps!